OpenStudy (anonymous):

s=1/2gt^2 solve for t

OpenStudy (e.mccormick):

\(s=\frac{1}{2}gt^2\) ?

OpenStudy (e.mccormick):

Solve for <variable> questions are basically order of operations problems, but you are using the order to un-do operations. Take \(x=\frac{y}{7}+1\) for example. The addition happens last in the normal order of operations. But if I say solve for y, then I need to un-do the addition first! This takes subtraction. \(x=\frac{y}{7}+1\) \(x-1=\frac{y}{7}+1-1\) \(x-1=\frac{y}{7}\) Then I need to un-do the division with multiplication. \(x-1=\frac{y}{7}\) \((x-1)\cdot 7=\frac{y}{7}\cdot 7\) \(7x-7=y\) Finally I rearrange it to a more standard order: \(y=7x-7\) That is all that this sort of question takes. So if you know your order of operations, just work it from the bottom to the top!